OpenStudy (anonymous):

I REALLY NEED HELP FAST! I WILL BECOME YOUR FAN/GIVE YOU A MEDAL IF U HELP ME?!?!! Select the equations that are parallel and perpendicular to y = x + 5 and that pass through the point (-2, -1

OpenStudy (anonymous):

parallel: y = -x - 1 perpendicular: y = x + 2 parallel: y = x - 1 perpendicular: y = -x + 1 parallel: y = x + 1 perpendicular: y = -x - 3 parallel: y = 2x - 2 perpendicular: y = -2x - 1

OpenStudy (anonymous):

@robtobey @dpaInc

OpenStudy (anonymous):

I don't know how to do this

OpenStudy (anonymous):

@radar @ganeshie8

OpenStudy (radar):

You can do this provided you understand the relationship between slopes of parallel vs perpendicular lines. Parallel lines will have the same slope (slopes are equal) Do you want to do the parallel line first?

OpenStudy (anonymous):

Yes please!

OpenStudy (anonymous):

Please hurry! I don't have much time cuz I have to go somewhere soon

OpenStudy (radar):

O.K. we want to calculate a parallel line to y=x+5 and it goes through the point (-2,-1) (x=-2, y=-1) The slope is the coefficient of the x in the y=x+5 equation so it is 1 (positive 1). Slope is not more than the change in y divided by the change in x. Or slope = (y1-y2)/(x1-x2) For this equation 1 = (y-(-1))/(x-(-2) or Did you follow ???